It took jose two thirds of an hour to complete his math homework on Monday,three -fourths of an hour on Tuesday and two-fifths of an hour on Wednesday , how many hours did it take jose to complete his homework altogether?
1
Expert's answer
2020-10-15T19:54:37-0400

2/3 of an hour -How much time Jose spend on hw on Monday

3/4 of an hour -How much time Jose spend on hw on Tuesday

2/5 of an hour -How much time Jose spend on hw on Wednesday

(2/3)+(3/4)+(2/5)=(2/3)*(20/20)+(3/4)*(15/15)+(2/5)*(12/12)=(40+45+24)/60=109/60(hours) or 109 minutes Jose spend to complete his homework altogether